OpenStudy (anonymous):

consider F(x)=(1)/(1+x^4)dt. x is from 2 to x. Find F'(x) and F'(2)....? Help??

OpenStudy (experimentx):

?? is it integration??

OpenStudy (anonymous):

i think so??

OpenStudy (anonymous):

\[\int\limits_{2}^{x} 1/(1+x^4) \]

OpenStudy (anonymous):

find f'(x) and f'(2)...?

OpenStudy (experimentx):

let, \[ \int\limits_{2}^{x} \frac1{(1+x^4)} dx = F(x) - F(2)\] \[ f'(x) = F'(x)*x' - 0 = \frac{1}{1+x^2}\] You can find f'(2) yourslef